Benzene has a refractive index of 1.43. Suppose a light ray strikes its


surface at an angle of 40°, what is the directions of the reflected and the


refracted rays?

1
Expert's answer
2022-05-22T14:37:38-0400

The angle of reflection is equal to the angle of incidence, so

"\theta_{\rm refl}=\theta_{\rm inc}=40^\circ"

The Snell's law says

"\sin\theta_{\rm inc} n_1=\sin\theta_{\rm refr}n_2"

Hence, the angle of refraction

"\theta_{\rm refr}=\sin^{-1}(\sin\theta_{\rm inc} n_1/n_2)"

"\theta_{\rm refr}=\sin^{-1}(\sin40^\circ* 1.00/1.43)\\
=26.7^\circ"